OpenStudy (anonymous):

What equation is the inverse of y = 16x^2 + 1?

OpenStudy (jhannybean):

Are you looking to find the inverse function of the one given?

OpenStudy (danjs):

Switch y and x around and solve for y

OpenStudy (anonymous):

Yes, @Jhannybean.

OpenStudy (jhannybean):

Ok, then I would start by what @DanJS has recommended as well.

OpenStudy (zzr0ck3r):

This function is NOT one-to-one, so there will be no inverse "function"

OpenStudy (zzr0ck3r):

you can find two different equations to model the inverse image

OpenStudy (jhannybean):

Ohh, because the function does not pass the horizontal line test?

OpenStudy (zzr0ck3r):

correct

OpenStudy (danjs):

ah right,

OpenStudy (zzr0ck3r):

But the original post said nothing about a function, so you can find the inverse relation.

OpenStudy (zzr0ck3r):

which is what you will get if you keep going like you were before I rudely butted in.

OpenStudy (anonymous):

I got \[y=\pm \sqrt{x-1}\div16\]

OpenStudy (jhannybean):

And the drawing is not to scale by the way, it's just a rough sketch of what it looks like.

OpenStudy (zzr0ck3r):

the 16 should be on the inside \(y=\pm\sqrt{\frac{x-1}{16}}=\pm\frac{\sqrt{x-1}}{4}\)

OpenStudy (anonymous):

ok

OpenStudy (jhannybean):